• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 08:02:31 on October11,2009

    Entry number 294374

    keyword=Owl Shift Summary

    00:08 MCC sending tune beam.
    00:17 start low beam run 13154

    00:25 back to 100uA.
    00:25 production run 13155, odd that the cuts + good events numbers don't add to 100%.

    1:22 requested 15uA by closing the slits for Compton tests.
    1:22 run 13156 at 15uA

    2:02 requested 100uA again

    There were DAQ problems (number of coda events did not increment), did a kcoda and then startcoda and DAQ is taking data again.

    2:10 Production run 13159
    3:07 Production run 13160

    4:01 MCC took away beam for path length studies.

    4:07 Pedestal run 13161

    4:28 MCC is sending tune beam
    4:36 MCC will send CW

    4:36 Production run 13162. The beam for this run is very trippy.

    5:11 requested 98uA from MCC if it will cause the beam to trip less. They quickly put it back up to 100uA.

    5:33 Production run 13163. Beam still very trippy.

    6:28 MCC is taking beam away to do work on the injector.

    6:30 Pedestal run 13164.

    7:11 MCC is returning to 100uA beam.
    7:11 Production run 13165.